x²-4x+3x-2=28 how do I solve this?
Answer by Edwin McCravy(20059)

x²-4x+3x-2 = 28

Combine like terms -4x and +3x as -1x

x²-1x-2 = 28

Add -28 to both sides to get 0 on the right:

x²-1x-30 = 0

Factor the left side:

(x-6)(x+5) = 0

Use the zero-factor property:

x - 6 = 0;  x + 5 = 0
    x = 6;      x = -5

Two solutions: 6 and -5

Some teachers like for you to write the solution set
by the roster method in order smallest to largest.:
 
{-5, 6}

Edwin